OpenStudy (anonymous):

I WI GIVE A MEDAL!! HELP!! how do I prove this to be false? X^k+1 divided by X equals X^k

jimthompson5910 (jim_thompson5910):

Why do you think it's false?

OpenStudy (anonymous):

@jim_thompson5910 Is it true?

OpenStudy (anonymous):

It doesn't seem true

OpenStudy (anonymous):

so x = x^k right ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It's x^x+1

OpenStudy (anonymous):

Here is a pic

jimthompson5910 (jim_thompson5910):

Rule \[\Large \frac{x^a}{x^b} = x^{a-b}\] It is true because \[\Large \frac{x^{k+1}}{x}\] \[\Large \frac{x^{k+1}}{x^1}\] \[\Large x^{(k+1)-(1)} ... \text{Use the rule given above}\] \[\Large x^{k+1-1}\] \[\Large x^{k+0}\] \[\Large x^{k}\] So that is why \[\Large \frac{x^{k+1}}{x} = x^k\] (where x is nonzero)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I gave you a medal

jimthompson5910 (jim_thompson5910):

glad to be of help